OpenStudy (babyslapmafro):

Please help, I am asked to sketch the curve by eliminating the parameter and indicate the direction of increasing t for the following parametric equations. (click to see).

OpenStudy (babyslapmafro):

x=2cos(t) y=5sin(t) 0 (greater than or equal to ) t (greater than or equal to) 2pi How do I determine what values for t are appropriate in this case?

OpenStudy (reemii):

Where is the picture?

OpenStudy (babyslapmafro):

There is no picture, I have to sketch it.

OpenStudy (reemii):

oh ok lol.

OpenStudy (reemii):

You see, there is a \(\cos(t)\) in the first coordinate, there is a \(\sin(t)\) in the second coordinate. The idea of the exercise is to forget about the \(t\) for an instant and look at a relation between \(x\) and \(y\). Do you know the fundamental formula involving \(\cos\alpha\) and \(\sin\alpha\)?

OpenStudy (babyslapmafro):

I know of addition and double-angle formulas.

OpenStudy (babyslapmafro):

and half-angle formulas

OpenStudy (reemii):

No, you must write the formula \(\cos^2t + \sin^2t=1\) in terms of \(x\) and \(y\). After doing so you should recognize the equation of a known curve.

OpenStudy (reemii):

\(\cos(t)=\frac x2\) and \(\sin(t)=\frac x5\) right? Write the formula and you will obtain the equation of an ellipse. It is easy to sketch (choose a few values for \(t\) to obtain some points you are sure of). To know the direction, choose \(t=0\) and then \(t=\pi/2\) for example, and look the direction in which the point has "moved".

OpenStudy (reemii):

I mean, \(\sin(t)=\frac y5\)

OpenStudy (jdoe0001):

$$ x=2cos(T) \implies cos(T)=\cfrac{{(x)}}{2}\\ y=5sin(T) \implies sin(T)=\cfrac{{(y)}}{5}\\ \text{now keeping in mind the pythagorean identity } cos(\theta)^2+sin(\theta)^2=1\\ \cfrac{(x)^2}{2^2}+\cfrac{(y)^2}{5^2}=1 \implies \cfrac{(x-0)^2}{2^2}+\cfrac{(y-0)^2}{5^2}=1 $$